OpenStudy (anonymous):

express using variables please 64^(1/6)

OpenStudy (anonymous):

\[\sqrt[6]{64}\]

OpenStudy (anonymous):

using variables? well i know 64=2^6 so technically your problem is just 2 and i don't know what you mean by express it in variables

OpenStudy (anonymous):

I'm not too sure what you mean. Something like x^6 = 64, so x=64^(1/6)=\[\pm2\]?

OpenStudy (anonymous):

\[x^2=2^6\] answer is 2 just guesin

jhonyy9 (jhonyy9):

yes 2 is one solution but the -2 is solution too --- check it please